OpenStudy (anonymous):

I CANT GET THIS!!!! PLEASE HELP What is the domain of the function y=1(over)x−1−1(over)x−2? A.{x | x ≥ 2} B.{x | x ≠ 1 and x ≠ 2} C.{x | x ≤ 1} d.R

OpenStudy (anonymous):

do you get this?

OpenStudy (anonymous):

yep.

OpenStudy (anonymous):

what u mean by 1 over x

OpenStudy (anonymous):

its a fraction 1/x-1 - 1/x-2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

its b

OpenStudy (anonymous):

wow thanks...! thanks I have 2 more questions do u think u can help me?

OpenStudy (anonymous):

i can try bt u got the answer of previous question?

OpenStudy (anonymous):

yes your very smart !... :)

OpenStudy (anonymous):

lol.i will not mind if u will give medals :P

OpenStudy (anonymous):

thankx:)

OpenStudy (anonymous):

Which equation does not define a function? A.y2 + x = 6 B.y + x3 = 3 C.y − x2 = 4 D.y3 − x2 = 7

OpenStudy (anonymous):

its a

OpenStudy (anonymous):

For the function f(x) = x2 − 9, what are a and b (the two values for f(x) missing from the table)?  a = 10 and b = -5 a = -8 and b = 5 a = 10 and b = 5 a = -8 and b = -5